[Brahms-l] Shift report 20050125 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Jan 25 2005 - 16:13:58 EST
1.Much trouble with the beam.
2.Got 2 data runs in the early morning (run 12481 & 12482). Did the pedestals after the beam was gone.
3.The magnets had some problems, but MCR sent people to fix it.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jan 25 16:14:16 2005

This archive was generated by hypermail 2.1.8 : Tue Jan 25 2005 - 16:14:39 EST